OpenStudy (makiadavis):

5x + 2y = 8 x + y = 4 If you want to solve the system of equations by addition, which of the following could you do? Multiply the second equation by 5 and add. Multiply the second equation by 2 and add. Multiply the second equation by -2 and add.

OpenStudy (jonathan34):

ok

OpenStudy (jonathan34):

no picture

OpenStudy (will.h):

what do you think? here they are asking you to solve using elimination method

OpenStudy (makiadavis):

no

OpenStudy (makiadavis):

yes

OpenStudy (the.lit.owl_23):

Multiply the second equation by -2 and add.

OpenStudy (ajspeller):

You need to eliminate either the x or y variable. Follow each of the given scenarios and whichever method causes the x or y to be eliminated would be the correct answer.

OpenStudy (the.lit.owl_23):

You could also multiply the second equation by -5 and add , "but that's not one of the choices "

OpenStudy (anglina4196):

5x+2y =8 x+y=4 Step 1) Multiply the 2nd equation by -5, so one of the variables cancel out 5x+2y=8 → 5x+2y=8 -5(x+y=4) → -5x-5y=-20 Step 2) Then add together the equations. The X's should cancel out. -3y=-12 Step 3) Then solve the equation -3y=-12 --- ---- -3 -3 y=4 That should be your y answer and to find the x just place in the 4 in the y of the equation. Just tell me if you need x